A sample of aluminium has a mass of 54 grand . What is the mass of the same number of magnesium atoms

Grade:11

1 Answers

Khimraj
6 years ago
Mole of aluminium = 54/27 = 2 mole. And atoms are 2Na where Na is Avogadro no. Then mole of magnesium is 2Na/Na = 2 and mass is 3*24 = 48 gran.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer